### v1

# Thermodynamic coprocessor for linear operations with input-size-independent calculation time based on open quantum system

Abstract
Linear operations, e.g., vector-matrix or vector-vector multiplications, are core operations of modern neural networks. To diminish computational time, these operations are implemented by parallel computations using different coprocessors. In this work we show that open quantum system consisting of bosonic modes and interacting with bosonic reservoirs can be used as analog coprocessor implementing multiple vector-matrix multiplications with stochastic matrices in parallel. Input vectors are encoded in occupancies of reservoirs, and output result is presented by stationary energy flows. The operation takes time needed for the system’s transition to non-equilibrium stationary state independently on number of the reservoirs, i.e., on the input vector dimension. The computations are accompanied by entropy growth. We construct a direct mapping between open quantum systems and electrical crossbar structures, showing that dissipation rates multiplied by OQS’s modes frequencies can be seen as conductivities, reservoirs’ occupancies can be seen as potentials, and stationary energy flows can be seen as electric currents.
preprint: APS/123-QED
Vector-matrix multiplication, particularly with stochastic matrices, is permanently exploited nowadays for natural language processing and steganography [1, 2, 3, 4], decision making [5, 6, 7], queuing theory [8], solving optimization problems such as the traveling salesman problem [9, 10, 11] and traffic optimization [12, 13, 14, 15], and for different types of forecasting, e.g., weather [16, 17], economy [18, 19, 20, 5], production [21, 22, 23, 24], social opinion [25, 26, 27]. In large-scale problems this operation is done by digital or analog coprocessors that implement it faster than a central processing unit due to parallel computations [28, 29, 30] or inherent physical features of the computational process [31]. Some of these platforms overcome the von Neumann bottleneck by in-memory computations  [32, 33, 34, 35, 36, 37] employing a crossbar structure (CS) for vector-matrix multiplication [38, 39, 40, 41, 42, 43, 44]. For instance, CSs are used in memristor-based neural networks (MNNs)  [45, 37, 35, 36] allowing MNNs to demonstrate low energy consumption and high suitability for neuromorphic computations [32, 33, 34, 35, 37, 36, 45, 46, 47].
A planar CS is a set of parallel conducting bars connected by memory (computing) elements with a similar one set of perpendicularly oriented bars [40, 42]. In MNN, semiconducting memristors are used for computations [37, 33, 32, 48, 45]. Usage of the memristors performing at quasi-particles [49, 50, 51, 52, 53, 54] and transition to quasi-particle MNN (QPMNN) can increase speed and efficiency of the network [53, 54, 55, 56]. For that, a proper CS should be designed to transmit the quasi-particles. With miniaturization, this technology will face quantum limit [57, 58, 59]. Thus, the design of a quantum CS, i.e., vector-matrix multiplayer, is a state-of-art problem.
The problem of the quantum CS design is very similar to the problems of quantum transport [60, 61, 62, 63] and the physics of open quantum systems (OQSs)  [64, 65, 66, 67]. In OQS’ physics, the non-Hermitian dynamics of a quantum system, i.e., OQS, connected to reservoirs (environments, i.e., sources of thermal noise), including dynamics of energy, particle, and heat flows [64, 68, 69, 70, 65, 71] is investigated [64, 72, 73, 74, 66, 67]. Reservoirs are considered large enough compared to the OQS. This allows to exclude their dynamics from the whole system’s quantum dynamics by Born-Markov approximation [64, 72]. After that, energy, particle, and heat flows between reservoirs connected via OQS can be calculated from the OQS’s dynamics [71, 75, 76]. During this evolution, entropy is growing [76] and OQS reaches its non-equilibrium stationary state [66], and stationary energy, particle, and heat flows from reservoirs are established [70, 65, 71]. Recently, this dynamics of OQS was proposed to implement matrix inversion and linear system of equations solution [77, 31]. Also, qubit-based OQS in Markovian limit was proposed as a adder [78, 79].
In this letter, we show that OQS consisting of bosonic modes can be used as an analog coprocessor implementing multiple vector-matrix multiplications with stochastic matrices in a parallel analog way. Input vectors are encoded in reservoirs’ occupancies, and output results are encoded in stationary energy flows. This takes the time needed for the transition of the OQS to its non-equilibrium stationary state. This time does not depend on the number of reservoirs. The consideration is done by means of global approach to dissipation [80, 71, 81], i.e., the second law of thermodynamics is fulfilled and computations are accompanied with entropy growth. We develop an electrical analogy for the OQS, showing that it can be represented as a planar CS circuit, i.e., dissipation rates multiplied by OQS’s modes frequencies can be seen as conductivities, reservoirs’ occupancies as potentials, and stationary energy flows as electric currents.
Let consider an OQS consisting of $K$ bosonic modes, see Fig. 1, that is a common model for photons [82, 83, 84, 85, 86], phonons [87, 88, 89], and magnons [90, 91, 92]. The Hamiltonian of this OQS reads [82, 83, 93] $\hat{H}_{S}=\sum_{\kappa=1}^{K}\omega_{\kappa}\hat{a}^{\dagger}_{\kappa}\hat{a}_{\kappa}.$ Here, $\omega_{\kappa}$-s are frequencies of the modes, $\hat{a}_{\kappa}$-s are lowering operators of the modes, $\hat{a}_{v},\hat{a}_{w}]=0$, $\hat{a}^{\dagger}_{v},\hat{a}_{w}]=-\delta_{vw}$. The number $K$ is arbitrary, but further we show that it defines the number of possible parallel vector-matrix multiplication operations.
Figure 1: Schematic representation of the considering OQS.
There are many approaches to describe the dynamics of the OQS interacting with a number of reservoirs [94, 66, 64]. In Markovian limit, local and global approaches are usually used [71, 81, 95, 70, 96]. However, under the local approach second law of thermodynamics can be violated [68, 70], while under the global approach it is always fulfilled [97]. With aim to ensure the fulfillment of second law of thermodynamics, further, we use the global approach.
Let OQS interacts with $n+1$ reservoirs at temperatures $T_{j}$, where $j$ goes from 0 to $n$. Further we suppose that $T_{0}\ll T_{j\neq 0}$ and use this cold reservoir as a drain. The energy flow from the $j$-th reservoir to the OQS in the global approach  [64, 98, 99] equals [65]
${J}_{j}=\sum_{\kappa=1}^{K}{J}_{\kappa,j}=\sum_{\kappa=1}^{K}\omega_{\kappa}\gamma_{\kappa,j}(n_{j}(\omega_{\kappa},T_{j})-\tilde{n}(\omega_{\kappa},T_{0,\cdots,n})).$ (1)
Here $\gamma_{\kappa,j}$ is dissipation rate of the OQS’s $\kappa$ mode (with frequency $\omega_{\kappa}$) to the $j$-th reservoir, $\tilde{n}(\omega_{\kappa},T_{0,\cdots,n})=\sum_{j=0}^{n}p_{\kappa,j}n_{j}(\omega_{\kappa},T_{j})$ is weighted occupancy among reservoirs at frequency $\omega_{\kappa}$, i.e. $p_{\kappa,j}=\gamma_{\kappa,j}/\sum_{m=0}^{n}\gamma_{\kappa,m}$, $n_{j}(\omega_{\kappa},T_{j})=(\exp(\omega_{\kappa}/T_{j})-1)^{-1}$ is occupancy of the $j$-th reservoir at frequency $\omega_{\kappa}$, ${J}_{\kappa,j}$ denotes energy flow through the frequency $\omega_{\kappa}$.
Let suppose that we can manage dissipation rates of OQS at different frequencies independently (there are a number of possibilities to do that even in the case of degenerate modes of OQS [100, 101, 102, 103, 104, 105, 106, 107, 108, 109]). Then, the dependence of energy flows in Eq. (1) on weighted occupancy $\tilde{n}(\omega_{\kappa},T_{0,\cdots,n})$ can be used to implement linear algebra operations as discussed below.
Scalar product. Eq. (1) can be rewritten as follows
${J}_{j}=\sum_{\kappa=1}^{K}{J}_{\kappa,j}=$ (2)
$=\sum_{\kappa=1}^{K}\omega_{\kappa}\sum\limits_{q=0}^{n}\frac{\gamma_{\kappa,j}\gamma_{\kappa,q}}{\sum\limits_{m=0}^{n}\gamma_{\kappa,m}}(n_{j}(\omega_{\kappa},T_{j})-n_{q}(\omega_{\kappa},T_{q})).$
Let $T_{0}\ll T_{j\neq 0}$ such that $n_{0}(\omega_{\kappa},T_{0})\ll n_{j\neq 0}(\omega_{\kappa},T_{j})$. Then,
${J}_{0}\approx-\sum_{\kappa=1}^{K}\omega_{\kappa}\gamma_{\kappa,0}\sum\limits_{q=1}^{n}p_{\kappa,q}n_{q}(\omega_{\kappa},T_{q})=$ (3)
$=-\sum_{\kappa=1}^{K}\omega_{\kappa}\gamma_{0,\kappa}(\vec{p}_{\kappa},\vec{n}(\omega_{\kappa},\vec{T})),$
$\vec{p}_{\kappa}=\left[\begin{array}[]{cc}p_{\kappa,1}\\ \vdots\\ p_{\kappa,n}\end{array}\right],\ \ \ \vec{n}(\omega_{\kappa},\vec{T})=\left[\begin{array}[]{cc}n_{1}(\omega_{\kappa},T_{1})\\ \vdots\\ n_{n}(\omega_{\kappa},T_{n})\end{array}\right].$ (10)
Here $\vec{T}=(T_{1},\ldots,T_{n})^{T}$.
It is seen that energy flow through $\omega_{\kappa}$ mode is proportional to scalar product of $\vec{p}_{\kappa}$ and $\vec{n}(\omega_{\kappa},\vec{T})$ vectors. This can be used to implement scalar product of positive vectors. Indeed, suppose that we want to find $(\vec{a},\vec{b})$, where $\vec{a},\vec{b}\in\mathds{R}^{n}\geq 0$. This denotes that elements of both vectors are non-negative. Vector $\vec{a}$ is considered to be normalized. For this purpose
(i) we need to connect $n+1$ reservoirs through a set of bosonic modes, one of the reservoirs should be cooled down $T_{0}\ll T_{j\neq 0}$ in such a way that for some $\omega_{\kappa}$ modes of the OQS holds $n_{0}(\omega_{\kappa},T_{0})\ll n_{j\neq 0}(\omega_{\kappa},T_{j})$ (the drain reservoir);
(ii) we need to choose a mode from them to conduct calculations, let it be $\omega_{1}$ mode;
(iii) we need to manipulate dissipation rates at frequency $\omega_{1}$ setting $\gamma_{0,1}\ll\sum_{j=1}^{n}\gamma_{j,1}$ to make $\vec{p}_{1}$ a normalized vector.
(iv) we need to manage dissipation rates at frequency $\omega_{1}$ and temperatures of other reservoirs to satisfy the relations
$\vec{p}_{1}=\vec{a},\quad\vec{n}(\omega_{1},\vec{T})=\vec{b}.$ (11)
(v) We need to analyze the spectrum of the energy outcome from the OQS to the drain reservoir and apply equation $(\vec{a},\vec{b})=-{J}_{1,0}/\omega_{1}\gamma_{1,0}$ that follows from Eq. (3).
For this procedure to be implemented, we need to set $\vec{b}$ by managing reservoirs’ temperatures. This can always be done as $\partial n(\omega,T)/\partial T>0$. Indeed, from $b_{j}=1/({\rm exp}(\omega_{1}/T_{j})-1)$ we get $T_{j}=\omega_{1}/\ln{(1+1/b_{j})}$.
Note that mode of the OQS in this setting acts as an adder. It sums its entries (reservoirs’ occupancies) with given weights (dissipative rates) and provides the output signal (energy flow) that is proportional to the weighted sum of the entries.
In parallel to the described procedure, different scalar products are calculated at other frequencies $\omega_{\kappa},\ \kappa\neq 1$. Indeed, as Eq. (3) states, the energy flows at other frequencies of OQS are also proportional to some scalar products. Occupancies of reservoirs at these frequencies are functionally dependent on occupancies at $\omega_{1}$. Hence, we get additional K1K-1 scalar products of vectors that are functionally dependent on $\vec{b}$ with different vectors defined by dissipation rates of OQS at frequencies $\omega_{\kappa},\ \kappa\neq 1$. This might be useful for feature extraction procedures in neural networks and object recognition [110, 111, 112].
Vector-matrix multiplication. Being able to implement the scalar product of a normalized positive vector with another positive vector, one can implement multiplication of an arbitrary stochastic matrix to a positive vector. For that we need to implement the scalar product several times by other modes of the OQS. To make the input vector $\vec{n}(\omega_{\kappa},\vec{T})$ approximately the same we need to use modes of the OQS with close frequencies.
Let the modes of OQS are split into $M$ groups with close frequencies ($M\leq K$), i.e. inside the group $n_{j\neq 0}(\omega_{\kappa},\vec{T})\approx n_{j\neq 0}(\omega_{1},\vec{T})$, and groups are well separated one from the other in frequency domaine. Let denote by $m_{1}$ the number of OQS’s modes that are close to $\omega_{1}$ such that $n_{j\neq 0}(\omega_{\kappa},\vec{T})\approx n_{j\neq 0}(\omega_{1},\vec{T})$. Applying the procedure described above to all these modes, we get
$\vec{{J}}_{1,0}=P_{1}\vec{n}(\omega_{1},\vec{T}),\ \text{where}$ (12)
$\vec{J}_{1,0}=\left[\begin{array}[]{cc}{J}_{1,0}\\ \vdots\\ {J}_{m_{1},0}\end{array}\right],\quad P_{1}=\left[\begin{array}[]{ccc}p_{1,1}&\cdots&p_{1,n}\\ \vdots&\ddots&\vdots\\ p_{m_{1},1}&\cdots&p_{m_{1},n}\end{array}\right].$ (19)
Every row of the matrix $P_{1}$ is a normalized vector. Hence, if $n=m_{1}$, then $P_{1}$ is a stochastic-row matrix, i.e. transition matrix of a Markov chain [8].
Note that in parallel with this procedure, vector-matrix multiplication happens at other close groups of $\omega_{\kappa}$, while different groups are well separated. This is similar to the way the parallel scalar product is implemented by different $\omega_{\kappa}$ described above. Input vectors $\vec{n}(\omega_{\kappa},\vec{T})$ of different groups are functionally dependent on the first $\vec{n}(\omega_{1},\vec{T})$ input vector. Thus, the results of these parallel computations are not totally independent but might be useful for feature extraction and object recognition [110, 111, 112].
The ability to manage dissipation rates can be used to implement learning process in this system. Indeed, with a feedback loop being designed, comparison of the computation result with the target result can be utilized for the dissipation rates’ managing [113, 112].
Important properties of the computational operations are their implementation time and fault-tolerance. The presented scheme uses stationary energy flows from noisy environments for computations. The establishment of the flows requires time that is of the order of the OQS dissipation time and does not depend on the initial state of the OQS and the number of reservoirs [65, 81], if long-living modes (for example, subradiant modes) are avoided in the evolution of the OQS [114, 115, 116, 117]. If, e.g., we consider an ensemble of resonators in the GHz range as the OQS, each of them interacts with phonons and free space electromagnetic modes as the drain reservoir. Then, for $\lambda\sim 10^{-3}$ m and Q-factor $0^{2}\div 10^{4} [11$8, 119], we can estimate the needed time as $10^{-8}\div 10^{-10}$ s. Notably, the lower Q-factor is, the faster computations are done.
In fact, these stationary energy flows are mean energy flows, as the fluctuations of temperatures are already averaged when Born-Markov approximation is valid [72, 120, 64]. This delivers reliable results due to the fact that reservoirs’ correlation functions decay faster than the OQS dissipates [72, 120, 64, 121, 65, 96]. Also, the OQS’s evolution to the stationary state is accompanied by entropy growth [64, 122, 97]. However, the OQS’s entropy can increase or decrees in this process. Nevertheless, this does not affect the result of the calculations. Mentioned effects mask the contribution of random errors and increase the fault-tolerance of the scheme.
Electrical analogy for OQS. Here we develop an electrical analogy for the considered OQS by mapping it to a CS. We show that the energy flow in Eq. (1) can be associated with electric current in electric circuit presented in Fig. 2. In this scheme $n$ wires having resistances $R_{j}$ and potentials $\varphi_{j}$ applied to their ends are connected by a central node. Let denote the potential of the central node $\tilde{\varphi}$.
The electric currents $I_{j}$ by Ohm’s law equal [123] $\varphi_{j}-\tilde{\varphi}=I_{j}R_{j}$. By Kirchhoff’s first law, the sum of incoming electric currents in the node should be zero in stationary regime $\sum_{j}I_{j}=0$ [123]. Hence, $\sum_{j}(\varphi_{j}-\tilde{\varphi})/R_{j}=0$ and
$\tilde{\varphi}=\frac{\sum_{j}\varphi_{j}/R_{j}}{\sum_{j}{1}/{R_{j}}}=\sum_{j}p_{j}\varphi_{j}.$ (20)
Here, $p_{j}=(1/R_{j})/\sum_{j}(1/R_{j})$. Thus, $\tilde{\varphi}$ equals the weighted potential of wires, with weights $p_{j}$ and
[…]
Figure 2: Electrical circuit supporting currents equivalent to the energy flow ${J}_{\kappa,j}(\vec{T})$ in Eq. (1) for a particular $\kappa$.
Comparing Eq. (21) with Eq. (1), one can easily see that they have similar linear structure. Indeed, for a particular $\kappa$ in Eq. (1), we can denote $1/R_{j}\rightarrow\omega_{\kappa}\gamma_{\kappa,j}$, $\varphi_{j}\rightarrow n_{j}(\omega_{\kappa},T_{j})$, $\tilde{\varphi}\rightarrow\tilde{n}(\omega_{\kappa},\vec{T})$ in Eq. (21). By that we instantly get ${J}_{\kappa,j}$ from $I_{j}$. Thus, dissipation rates of OQS multiplied by modes’ frequencies play the role of resistances at power minus one (conductivities), and occupancies of OQS’s reservoirs are equivalent to potentials.
As there are many $\omega_{\kappa}$ modes in the OQS, the total system can be considered as a set of circuits from Fig. 2. Transforming the central node into a bar with no resistance, we get the circuit in Fig. 3. Here we have denoted $R_{\kappa,j}\equiv 1/\omega_{\kappa}\gamma_{\kappa,j},\ \varphi_{\kappa,j}\equiv n_{j}(\omega_{\kappa},T_{j})$ (where $\varphi_{\kappa,0}=0$).
If we connect all $\varphi_{\kappa,j}$ with the same $j$ through resistors to provide the decrease in potential, we get a crossbar representation of the considered OQS, see Fig. 4. Here
$\Phi_{j}(T_{j})\equiv\varphi_{\kappa,j}+I_{\kappa,j}(\vec{T})r_{\kappa,j}(\vec{T}),$ (22)
currents $I_{\kappa,j}(\vec{T})\equiv{J}_{\kappa,j}(\vec{T})/\omega_{\kappa}$. Note, the $\Phi_{j}(T_{j})$ is free from index $\kappa$, and depends only on $T_{j}$.
From Eqs. (22) it is seen that $\Phi_{j}(T_{j})$ depend on $r_{\kappa,j}(\vec{T})$, while $r_{\kappa,j}(\vec{T})$ are free parameters. This set of equations always has a solution. Indeed, let potentials $\Phi_{j}(T_{j})=\max_{\kappa}n_{j}(\omega_{\kappa},T_{j})$. Then the resistances required for the decrease in potentials equal $r_{\kappa,j}(\vec{T})=(\Phi_{j}(T_{j})-n_{j}(\omega_{\kappa},T_{j}))/I_{\kappa,j}(\vec{T})$. More over, this is not the unique solution of Eq. (22). For $\Phi_{j}$ such as $\Phi_{j}>\max_{\kappa}n_{j}(\omega_{\kappa},T_{j})$, there always exists a set of resistors $r_{\kappa,j}$ that satisfies Eqs. (22). Hence, there are infinitely many sets of $\Phi_{j}$ and $r_{\kappa,j}$ values that satisfy Eqs. (22).
Figure 3: Electrical circuit supporting currents equivalent to the energy flows ${J}_{\kappa,j}(\vec{T})$ in Eq. (1).
Figure 4: Equivalent CS of the OQS. Potentials $\varphi_{\kappa,}$ from Fig. 3 with the same $j$ are connected by common bars through resistors $r_{\kappa,j}$. Potentials $\Phi_{j}$ from Eqs. (22) are applied to these bars, providing the same currents as in Fig. 3.
For the considered computational scheme of vector-matrix multiplication based on OQS in each group of close modes we have $\Phi_{j}(T_{j})=n_{j}(\omega_{\kappa}/T_{j})$ and $r_{\kappa,j}=0$. Also, $\varphi_{0}=0$.
As it is shown, electric currents in the circuit presented in Fig. 4 are equivalent to stationary energy flows in the OQS of bosonic modes connected to several reservoirs with different temperatures. The dissipation rates in this OQS are equivalent to resistances at power minus one (conductivities), while reservoirs’ occupancies and temperatures can be seen as potentials (see Figs. 2-4).
By this analogy, we can expand the class of matrices that can be used in computations similarly to how it is done in MNN. If we need to multiply a matrix $A$ whose rows are normalized vectors with positive and negative elements by a positive vector $\vec{b}$, we can split it into two matrices $A=A_{+}-A_{-}$. The $A_{+}$ cares all positive elements in their places in the matrix $A$ and zeros in other places, while the $A_{-}$ cares all absolute values of negative elements of $A$ in their places in $A$, and zeros in other places. After that, the procedure using Eq. (12) can be applied to both matrices $A_{+}$ and $A_{-}$. The results being subtracted with attenuation deliver $A\vec{b}$.
To summarize, in this paper we show that OQS of bosonic modes, i.e., photons, phonons, magnons, etc., can be used as coprocessors implementing multiple vector-matrix multiplications in a parallel analog way with stochastic matrices. We show that stationary energy flow through one of the OQS’s modes to a reservoir sufficiently colder than others (drain reservoir) is proportional to the scalar product of vectors composed of reservoirs’ occupancies and weighted dissipation rates. We show that this fact and ability to manage dissipation rates of OQS’s modes with close frequencies allow us to implement multiplication of an $n$-dimensional vector by an $m\times n$-dimensional matrix, where $n+1$ is the number of reservoirs and $m$ is the number of the modes having close frequencies. The procedure requires the time needed for the establishment of the stationary energy flows, i.e., the OQS dissipation time, which does not depend on the number of reservoirs (input vector dimension) and the initial state of the OQS. In parallel, vector-matrix multiplications with different matrices are implemented at other close frequency groups of OQS’s modes.
The evolution of OQS under interaction with thermal environments is treated in global approach to dissipation [80, 71, 81] that ensures the fulfillment second law of thermodynamics. Hence, the computations being implemented by energy flows are accompanied by entropy growth due to OQS evolution towards its stationary state. Thus, the presented procedure can be seen as a part of developing fields of thermal linear algebra, thermodynamic computing, and thermal neural networks [77, 31, 78, 79].
Notably, intense interaction with the environment in many applications is considered to be a drawback, i.e., entanglement, coherence, and energy should be preserved as long as possible in the OQS  [124, 95, 125]. In the considered procedure the situation is opposite. The faster OQS dissipates, the faster computations are done.
We develop an electrical analogy for the considered system. We show that the system is equivalent to a CS. The dissipation rates of the OQS multiplied by OQS’s modes’ frequencies can be seen as conductivities, reservoirs’ occupancies as potentials, and stationary energy flows as electric currents.
Also we note, that the similar dependence of energy and particle flows on the weighted occupancy of reservoirs is found in fermionic systems with quadratic Hamiltonians [76], i.e. the same electrical analogy holds true in this case and these systems can also be used for the implementation of vector-matrix multiplication at the nanoscale.
Acknowledgement. A.A.Z. and E.S.A. acknowledge the support of the Foundation for the Advancement of Theoretical Physics and Mathematics BASIS.