### v2

# Thermodynamic coprocessor for linear operations with input-size-independent calculation time based on open quantum system

Abstract
Linear operations, e.g., vector-matrix and vector-vector multiplications, are core operations of modern neural networks. To diminish computational time, these operations are implemented by parallel computations using different coprocessors. In this work we show that an open quantum system consisting of bosonic modes and interacting with bosonic reservoirs can be used as an analog thermodynamic coprocessor implementing multiple vector-matrix multiplications with stochastic matrices in parallel. Input vectors are encoded in occupancies of reservoirs, and the output result is presented by stationary energy flows. The operation takes time needed for the system’s transition to a non-equilibrium stationary state independently on the number of the reservoirs, i.e., on the input vector dimension. With technological limitations being considered, a device of $5\times 5$ cm2 area covered with the coprocessors can conduct of the order of $10^{11}$ operations per second per a mode of the OQS. The computations are accompanied by an entropy growth. We construct a direct mapping between open quantum systems and electrical crossbar structures frequently used in analog vector-matrix multiplication, showing that dissipation rates multiplied by open quantum system’s modes frequencies can be seen as conductivities, reservoirs’ occupancies can be seen as potentials, and stationary energy flows can be seen as electric currents.
preprint: APS/123-QED
Introduction. Vector-matrix multiplication, particularly with stochastic matrices, is permanently exploited nowadays for natural language processing and steganography [6, 143, 26, 111], decision making [10, 4, 7], queuing theory [15], solving optimization problems such as the traveling salesman problem [39, 82, 75], traffic optimization [103, 2, 136, 73], and for different types of forecasting, e.g., weather [137, 31], economy [92, 59, 83, 10], production [34, 29, 33, 32], social opinion [52, 9, 147]. In large-scale problems this operation is done by digital or analog coprocessors that implement it faster than a central processing unit (CPU) due to parallel computations [53, 12, 18]. Some of the used platforms overcome the von Neumann bottleneck by in-memory computations  [128, 107, 13, 116, 72, 28] employing a crossbar structure (CS) for analog vector-matrix multiplication [134, 22, 67, 55, 102, 24, 35, 145]. Particularly, the CSs are used in memristor-based neural networks (MNNs)  [1, 28, 116, 72] allowing MNNs to demonstrate low energy consumption and high suitability for neuromorphic computations [128, 107, 13, 116, 28, 72, 1, 115, 108].
A planar CS is a set of parallel conducting bars connected by memory (computing) elements with a similar set of perpendicularly oriented bars [67, 102]. In MNN, semiconducting memristors are used for computations [28, 107, 128, 81, 1]. Usage of the memristors performing at quasi-particles [113, 144, 25, 140, 74, 110] and transition to quasi-particle MNN (QPMNN) can increase speed and efficiency of the network [74, 110, 69, 20]. For that, a proper CS should be designed to transmit the quasi-particles. With miniaturization, this technology will face quantum limit [121, 133, 96]. Thus, the design of a quantum CS, i.e., vector-matrix multiplayer, is a state-of-art problem.
The problem of the quantum CS design is very similar to the problems of quantum transport [139, 58, 88, 45] and the physics of open quantum systems (OQSs) [17, 125, 14, 36]. In OQS’ physics, the non-Hermitian dynamics of a quantum system, i.e., OQS, connected to reservoirs (environments, i.e., sources of thermal noise), including dynamics of energy, particle, and heat flows [17, 66, 119, 124, 125, 93, 105, 91] is investigated [17, 19, 98, 122, 14, 36]. Reservoirs are considered large enough compared to the OQS. This allows to exclude their dynamics from the whole system’s quantum dynamics by Born-Markov approximation [17, 19]. After that, energy, particle, and heat flows between reservoirs connected via OQS can be calculated from the OQS’s dynamics [93, 109, 126]. During this dynamics, entropy is growing [126] and OQS reaches its non-equilibrium stationary state [14], and stationary energy, particle, and heat flows from reservoirs are established [124, 125, 93].
Recently, the transition of OQS to its stationary state was proposed to implement matrix inversion and to solve linear system of equations [3, 77]. Also, qubit-based OQS in Markovian limit was proposed as a adder [70, 118]. Such an involvement of the inherent thermodynamic physical features into the computational process is of interest nowadays. The thermodynamic approach to computations shows great speed and low energy consumption [77]. Particularly, this is important in the modern trend of rapidly growing data centers’ energy demand [51, 50].
In this letter, we show that OQS consisting of bosonic modes can be used as a thermodynamic analog coprocessor implementing multiple vector-matrix multiplications with stochastic matrices in a parallel analog way. Input vectors are encoded in reservoirs’ occupancies, and output results are encoded in stationary energy flows. This takes the time needed for the transition of the OQS to its non-equilibrium stationary state. This time does not depend on the number of reservoirs and afford $10\div 1000$ TOps/s (terraoperations per second) on the $5\times 5$ cm2 area per a mode of the OQS. However, nowadays technological opportunities in temperature manipulation reduce this computational rate to $100$ GOps/s per. The consideration of the OQS’s dynamics is done by means of the global approach to dissipation [47, 93, 127], i.e., the second law of thermodynamics is fulfilled and computations are accompanied with entropy growth. We develop an electrical analogy for the OQS, showing that it can be represented as a planar CS circuit, i.e., dissipation rates multiplied by OQS’s modes frequencies can be seen as conductivities, reservoirs’ occupancies as potentials, and stationary energy flows as electric currents.
The model. We consider an OQS consisting of $K$ bosonic modes, see Fig. 1, that is a common model for photons [43, 44, 129, 100, 101], phonons [80, 142, 106], and magnons [141, 56, 76]. The Hamiltonian of this OQS reads [43, 44, 8] $\hat{H}_{S}=\sum_{\kappa=1}^{K}\omega_{\kappa}\hat{a}^{\dagger}_{\kappa}\hat{a}_{\kappa}.$ Here, $\omega_{\kappa}$ are frequencies of the modes, $\hat{a}_{\kappa}$ are lowering operators of the modes, $\hat{a}_{v},\hat{a}_{w}]=0$, $\hat{a}^{\dagger}_{v},\hat{a}_{w}]=-\delta_{vw}$. The number $K$ is arbitrary, but further we show that it defines the number of possible parallel vector-matrix multiplication operations in the OQS.
Figure 1: Schematic representation of the considering OQS.
There are many approaches to describe the dynamics of the OQS interacting with a number of reservoirs [94, 14, 17]. In the Markovian limit, local and global approaches are usually used [93, 127, 123, 124, 21]. However, under the local approach, second law of thermodynamics can be violated [66, 124], while under the global approach it is always fulfilled [114]. With the aim to ensure the fulfillment of the second law of thermodynamics, we use the global approach.
We suppose that the OQS interacts with $n+1$ reservoirs at temperatures $T_{j}$, where $j$ goes from 0 to $n$. Further, we assume that $T_{0}\ll T_{j\neq 0}$ and use this cold reservoir as a drain. The energy flow from the $j$-th reservoir to the OQS in the global approach  [17, 38, 27] equals [125]
${J}_{j}=\sum_{\kappa=1}^{K}{J}_{\kappa,j}=\sum_{\kappa=1}^{K}\omega_{\kappa}\gamma_{\kappa,j}(n_{j}(\omega_{\kappa},T_{j})-\tilde{n}(\omega_{\kappa},T_{0,\cdots,n})).$ (1)
Here $\gamma_{\kappa,j}$ is dissipation rate of the OQS’s $\kappa$ mode (with frequency $\omega_{\kappa}$) to the $j$-th reservoir, $\tilde{n}(\omega_{\kappa},T_{0,\cdots,n})=\sum_{j=0}^{n}p_{\kappa,j}n_{j}(\omega_{\kappa},T_{j})$ is weighted occupancy among reservoirs at frequency $\omega_{\kappa}$, i.e., $p_{\kappa,j}=\gamma_{\kappa,j}/\sum_{m=0}^{n}\gamma_{\kappa,m}$, $n_{j}(\omega_{\kappa},T_{j})=(\exp(\omega_{\kappa}/T_{j})-1)^{-1}$ is occupancy of the $j$-th reservoir at frequency $\omega_{\kappa}$, ${J}_{\kappa,j}$ denotes energy flow through the frequency $\omega_{\kappa}$. Note that the $\omega_{\kappa}$ used here from [125] can be seen as eigenmodes’ frequencies of an diaganolized quadratic Hamiltonian that takes the coupling between modes into the account [11, 16]. Here and further, we address these eigenmodes as modes. This approach is reliable when nonlinearities in the OQS or its modes’ amplitudes are small [57, 37, 65]. Then the contribution of nonlinearities can also be taken into account by a diagonalizable quadratic Hamiltonian [57, 142, 64].
The dependence of energy flows in Eq. (1) on weighted occupancy $\tilde{n}(\omega_{\kappa},T_{0,\cdots,n})$ can be used to implement linear algebra operations as discussed below.
The scalar product. Eq. (1) can be rewritten as follows
${J}_{j}=\sum_{\kappa=1}^{K}{J}_{\kappa,j}=$ (2)
$=\sum_{\kappa=1}^{K}\omega_{\kappa}\sum\limits_{q=0}^{n}\frac{\gamma_{\kappa,j}\gamma_{\kappa,q}}{\sum\limits_{m=0}^{n}\gamma_{\kappa,m}}(n_{j}(\omega_{\kappa},T_{j})-n_{q}(\omega_{\kappa},T_{q})).$
We consider $T_{0}\ll T_{j\neq 0}$ such that $n_{0}(\omega_{\kappa},T_{0})\ll n_{j\neq 0}(\omega_{\kappa},T_{j})$. Then,
${J}_{0}\approx-\sum_{\kappa=1}^{K}\omega_{\kappa}\gamma_{\kappa,0}\sum\limits_{q=1}^{n}p_{\kappa,q}n_{q}(\omega_{\kappa},T_{q})=$ (3)
$=-\sum_{\kappa=1}^{K}\omega_{\kappa}\gamma_{0,\kappa}(\vec{p}_{\kappa},\vec{n}(\omega_{\kappa},\vec{T})),$
$\vec{p}_{\kappa}=\left[\begin{array}[]{cc}p_{\kappa,1}\\ \vdots\\ p_{\kappa,n}\end{array}\right],\ \ \ \vec{n}(\omega_{\kappa},\vec{T})=\left[\begin{array}[]{cc}n_{1}(\omega_{\kappa},T_{1})\\ \vdots\\ n_{n}(\omega_{\kappa},T_{n})\end{array}\right].$ (10)
Here $\vec{T}=(T_{1},\ldots,T_{n})^{T}$.
It is seen that energy flow through the $\omega_{\kappa}$ mode is proportional to the scalar product of $\vec{p}_{\kappa}$ and $\vec{n}(\omega_{\kappa},\vec{T})$ vectors. This can be used to implement scalar product of positive vectors. Indeed, suppose that we want to find $(\vec{a},\vec{b})$, where $\vec{a},\vec{b}\in\mathds{R}^{n}\geq 0$ (this denotes that elements of both vectors are non-negative). Vector $\vec{a}$ is considered to be normalized. For that
(i) one needs to connect $n+1$ reservoirs through a set of bosonic modes, and one of the reservoirs should be cooled down $T_{0}\ll T_{j\neq 0}$ to fulfill the condition $n_{0}(\omega_{\kappa},T_{0})\ll n_{j\neq 0}(\omega_{\kappa},T_{j})$ for some modes $\omega_{\kappa}$ of the OQS (the drain reservoir);
(ii) we need to choose a mode from the modes mentioned above to conduct the calculations, let it be the $\omega_{1}$ mode;
(iii) we need to manipulate dissipation rates at frequency $\omega_{1}$ setting $\gamma_{0,1}\ll\sum_{j=1}^{n}\gamma_{j,1}$ to make $\vec{p}_{1}$ a normalized vector.
(iv) we need to manage dissipation rates at frequency $\omega_{1}$ and temperatures of other reservoirs to satisfy the relations
$\vec{p}_{1}=\vec{a},\quad\vec{n}(\omega_{1},\vec{T})=\vec{b};$ (11)
(v) we need to analyze the spectrum of the energy outcome from the OQS to the drain reservoir and apply the equation $(\vec{a},\vec{b})=-{J}_{1,0}/\omega_{1}\gamma_{1,0}$ that follows from Eq. (3).
For this procedure to be implemented, one needs to set $\vec{b}$ by managing reservoirs’ temperatures. This can always be done as $\partial n(\omega,T)/\partial T>0$. Indeed, from $b_{j}=1/({\rm exp}(\omega_{1}/T_{j})-1)$ we get $T_{j}=\omega_{1}/\ln{(1+1/b_{j})}$. Management of the dissipation rates can be done in many ways, particularly by managing the coupling between the OQS’s modes and the modes of the reservoirs [95, 87, 78, 84, 85, 60, 61, 23, 149, 89].
Note that the mode of the OQS in this setting acts as an adder. It sums its entries (reservoirs’ occupancies) with given weights (dissipative rates) and provides the output signal (energy flow) that is proportional to the weighted sum of the entries. One can restrict the values of the entries at a moderate level of the OQS modes’ amplitudes to process normalized data. Then the contribution of OQS’s nonlinearities would be taken into account by the diagonalized OQS’s Hamiltonian, as it is discussed above.
In parallel to the described procedure, different scalar products are calculated at other frequencies $\omega_{\kappa},\ \kappa\neq 1$. Indeed, as Eq. (3) states, the energy flows at other frequencies of OQS are also proportional to some scalar products. Occupancies of reservoirs at these frequencies are functionally dependent on occupancies at $\omega_{1}$. Hence, we get additional K1K-1 scalar products of vectors that are functionally dependent on $\vec{b}$ with different vectors defined by dissipation rates of OQS at the frequencies $\omega_{\kappa},\ \kappa\neq 1$. This might be useful for feature extraction procedures in neural networks and object recognition [40, 86, 68].
For that, the frequency distance between the modes of the OQS should be sufficient to neglect the effect of non-resonant stationary energy transport, which is revealed in non-Markovian models of the OQS dynamics [45, 105, 91]. This effect is relevant when the frequency distance between OQS’s modes is of the order of the dissipation rates.
The vector-matrix multiplication. Being able to implement the scalar product of a normalized positive vector with another positive vector, one can implement multiplication of an arbitrary stochastic matrix to a positive vector. For that, we need to implement the scalar product several times. This can be done by employing an array of considered systems or by employing other modes of the OQS.
The first option is trivial. If one has $m$ copies of the system, then
$\vec{{J}}_{1,0}=P_{1}\vec{n}(\omega_{1},\vec{T}),\ \text{where}$ (12)
$\vec{J}_{1,0}=\left[\begin{array}[]{cc}{J}_{1,0}^{(1)}\\ \vdots\\ {J}_{1,0}^{(m)}\end{array}\right],\quad P_{1}=\left[\begin{array}[]{ccc}p_{1,1}^{(1)}&\cdots&p_{1,n}^{(1)}\\ \vdots&\ddots&\vdots\\ p_{1,1}^{(m)}&\cdots&p_{1,n}^{(m)}\end{array}\right].$ (19)
Here ${J}_{1,0}^{(k)}$ is energy flow through the $\omega_{1}$ mode of the $k$-th copy of the OQS, and $p_{1,1}^{(k)}$ is $p_{1,1}$ in the $k$-th copy of the OQS. Every row of the matrix $P_{1}$ is a normalized positive vector. Hence, if $n=m$, then $P_{1}$ is a stochastic-row matrix, i.e., transition matrix of a Markov chain [15].
Note that in parallel with this procedure, vector-matrix multiplication happens at other modes of the OQS. This is similar to the way the parallel scalar product is implemented by different $\omega_{\kappa}$ described above. Input vectors $\vec{n}(\omega_{\kappa},\vec{T})$ at different modes are functionally dependent on the first $\vec{n}(\omega_{1},\vec{T})$ input vector. Thus, the results of these parallel computations are not totally independent but might be useful for feature extraction and object recognition [40, 86, 68].
To implement the second option the input vectors $\vec{n}(\omega_{\kappa},\vec{T})$ should be approximately the same at the used OQS’s modes. Thus, we need to use degenerate modes of the OQS or modes with close frequencies. If one handles independent management of the dissipation rates of OQS’s modes with different frequencies, this second option can be realized multiple times in the frequency domain of the OQS with several well-separated-by-frequency-gaps groups of the modes. Along with the copying of the system, this can boost the total rate of computations per unit area.
The ability to manage dissipation rates can also be used to implement learning process in this system. Indeed, with a feedback loop being designed, comparison of the computation result with the target result can be utilized for the dissipation rates’ managing [30, 68].
The device. In this section we discuss possible realization of the proposed computational scheme. The important properties of a computational device are number of operations done per second and fault tolerance. The presented scheme uses stationary energy flows from noisy environments for computations. The establishment of the flows requires time that is of the order of the OQS dissipation time and does not depend on the initial state of the OQS and the number of the reservoirs [125, 127], if long-living modes (for example, subradiant modes) are avoided in the evolution of the OQS [132, 54, 148, 79]. If, e.g., we consider an ensemble of resonators in the GHz range as the OQS, interacting with the free space electromagnetic modes as the drain reservoir, then for the wavelength of the OQS mode $\sim 10^{-3}$ m and Q-factor (quality factor) $\sim 10^{2}\div 10^{4}$ that can be achieved simultaneously in the experiments with microrings and photonic crystal resonators [90, 62, 117, 48], we can estimate the computational time as $10^{-8}\div 10^{-10}$ s. This time is required to realize the above-discussed vector-matrix multiplication. Notably, the lower the Q-factor is, the faster computations are done.
To implement the reservoirs of the OQS, one can use locally heated waveguides [120, 146, 135]. The heating can be provided by heaters that are of $10^{-5}$ m size [120, 146, 135]. The dissipation rates then are proportional to the square of evanescent coupling between the waveguides’ modes and the modes of the OQS [125, 138]. This coupling can be controlled in multiple ways [95, 87, 78, 84, 85, 60, 61, 23, 149, 89, 138, 49, 46] (for example, via management of the distance between OQS’s elements and the waveguides). Then, one can expect the $n\sim 100$ on the $10^{-3}$ m size of the OQS mode’s wavelength and the computational rate of about $10\div 1000$ GOps/s per a mode of the OQS.
On the approximate CPU area of $5\times 5$ cm2, it is possible to place of the order of $1000$ of the considered computing systems. Hence, the possible computational rate of such a device per a mode of the OQS can be estimated as $10\div 1000$ TOps/s.
That large computational rate is a fundamental limitation of the computational rate per the limited area. To reach it, the whole computational cycle should be done on the timescale of the OQS dissipation time. Here, the technological limitations impose stricter restrictions. The computational rate of the coprocessor in reality will be limited by the time of thermal response that is usually on the order of microseconds in the experiments [120, 146]. Hence, the nowadays realistic value of the coprocessor computational rate should be estimated as $100$ GOps/s per the OQS’s mode.
This computational rate is comparable with the computational rate of some modern GPU solutions used in industry [97]. Furthermore, it is of great interest in the context of a full-thermodynamic computer development [3, 77, 70, 118, 99, 131, 130] and extra computations in microelectronic systems conducted by spurious heat [112]. Also, the development of faster heating elements along with the design of the scheme allowing for more guided and direct heating effect can increase the computational rate of the device.
Additionally, one can try managing dissipation rates of the OQS modes having different frequencies in parallel. As it is discussed above, this can increase the computational rate in the system too. For that, a proper choice of the waveguides and the medium that connect them to the OQS should be done to match the needed integral of overlapping between the electromagnetic fields of the OQS and the reservoirs at chosen frequencies [138, 49, 46].
Note that the stationary energy flows in OQS in fact are mean stationary energy flows, as the fluctuations of temperatures are already averaged when Born-Markov approximation is valid [19, 42, 17]. This delivers reliable results due to the fact that reservoirs’ correlation functions decay faster than the OQS dissipates [19, 42, 17, 41, 125, 21]. Also, the OQS’s evolution to the stationary state is accompanied by entropy growth [17, 63, 114]. However, the OQS’s entropy can increase or decrease in this process. Nevertheless, this does not affect the result of the calculations. Mentioned effects mask the contribution of random errors and increase the fault tolerance of the scheme.
Also, we should mention that the influence of the non-Markovian non-resonant stationary energy transport discussed above is negligible in the considering scheme. Indeed, the dissipation rate of the OQS is inversely proportional to its dissipation time and defines the spectral width of the non-Markovian non-resonant stationary energy transport [45, 105]. Thus, in the considering system, this spectral width is of the order of $10^{8}\div 10^{10}$ s-1, while the frequency step between OQS’s modes is of the order of $10^{11}$ s-1. Hence, the non-Markovian non-resonant stationary energy transport is irrelevant in the considering setup.
The discussed realization of the thermodynamic coprocessor has restricted opportunities for the dissipation rates manipulation. Thus, it is a worthwhile device when a multiplication of a particularly needed, once-established matrix by an arbitrary normalized positive vector should be done. For instance, the considered realization of the coprocessor is suitable for performance in an already learned neural network.
Electrical analogy for the OQS. As CS are usually used in microelectronics, it will be useful to have an electrical analogy of the process considered above to visualize the achieved results. Here we develop it by mapping the OQS to a CS. We show that the energy flow in Eq. (1) can be associated with electric current in the electric circuit presented in Fig. 2.
In this circuit $n$ wires having resistances $R_{j}$ and potentials $\varphi_{j}$ applied to their ends are connected by a central node. The potential formed in the central node is denoted as $\tilde{\varphi}$. The electric currents $I_{j}$ by Ohm’s law equal [5] $\varphi_{j}-\tilde{\varphi}=I_{j}R_{j}$. By Kirchhoff’s first law, the sum of incoming electric currents in the node should be zero in stationary regime $\sum_{j}I_{j}=0$ [5]. Hence, $\sum_{j}(\varphi_{j}-\tilde{\varphi})/R_{j}=0$ and
$\tilde{\varphi}=\frac{\sum_{j}\varphi_{j}/R_{j}}{\sum_{j}{1}/{R_{j}}}=\sum_{j}p_{j}\varphi_{j}.$ (20)
Here, $p_{j}=(1/R_{j})/\sum_{j}(1/R_{j})$. Thus, $\tilde{\varphi}$ equals the weighted potential of wires, with weights $p_{j}$ and
[…]
Figure 2: Electrical circuit supporting currents equivalent to the energy flow ${J}_{\kappa,j}(\vec{T})$ in Eq. (1) for a particular $\kappa$, i.e., the energy flows through a frequency of the OQS.
Comparing Eq. (21) with Eq. (1), one can easily see that they have similar linear structure. Indeed, for a particular $\kappa$ in Eq. (1), we can denote $1/R_{j}\rightarrow\omega_{\kappa}\gamma_{\kappa,j}$, $\varphi_{j}\rightarrow n_{j}(\omega_{\kappa},T_{j})$, $\tilde{\varphi}\rightarrow\tilde{n}(\omega_{\kappa},\vec{T})$ in Eq. (21). By that we instantly get ${J}_{\kappa,j}$ from $I_{j}$. Thus, dissipation rates of OQS multiplied by modes’ frequencies play the role of resistances at power minus one (conductivities), and occupancies of OQS’s reservoirs are equivalent to potentials.
As there are many $\omega_{\kappa}$ modes in the OQS, the total system can be considered as a set of circuits from Fig. 2. Transforming the central node into a bar with no resistance and repeating this circuit, we get the circuit in Fig. 3. Here we have denoted $R_{\kappa,j}\equiv 1/\omega_{\kappa}\gamma_{\kappa,j},\ \varphi_{\kappa,j}\equiv n_{j}(\omega_{\kappa},T_{j})$ (note, $\varphi_{\kappa,0}=0$).
If we connect all $\varphi_{\kappa,j}$ with the same $j$ through resistors to provide the decrease in potential, we get a crossbar representation of the considered OQS, see Fig. 4. Here
$\Phi_{j}(T_{j})\equiv\varphi_{\kappa,j}+I_{\kappa,j}(\vec{T})r_{\kappa,j}(\vec{T}),$ (22)
currents $I_{\kappa,j}(\vec{T})\equiv{J}_{\kappa,j}(\vec{T})/\omega_{\kappa}$. Note, the $\Phi_{j}(T_{j})$ is free from index $\kappa$ and depends only on $T_{j}$.
From Eqs. (22) it is seen that $\Phi_{j}(T_{j})$ depend on $r_{\kappa,j}(\vec{T})$, while $r_{\kappa,j}(\vec{T})$ are free parameters. This set of equations always has a solution. Indeed, if potentials $\Phi_{j}(T_{j})=\max_{\kappa}n_{j}(\omega_{\kappa},T_{j})$, then the resistances required for the decrease in potentials equal $r_{\kappa,j}(\vec{T})=(\Phi_{j}(T_{j})-n_{j}(\omega_{\kappa},T_{j}))/I_{\kappa,j}(\vec{T})$. More over, this is not the unique solution of Eq. (22). For $\Phi_{j}$ such that $\Phi_{j}>\max_{\kappa}n_{j}(\omega_{\kappa},T_{j})$, there always exists a set of resistors $r_{\kappa,j}$ that satisfies Eqs. (22). Hence, there are infinitely many sets of $\Phi_{j}$ and $r_{\kappa,j}$ values that satisfy Eqs. (22).
Figure 3: Electrical circuit supporting currents equivalent to the energy flows ${J}_{\kappa,j}(\vec{T})$ in Eq. (1). This circuit is the circuit from Fig. 2 repeated for each mode of the OQS.
Figure 4: Equivalent CS of the OQS. Potentials $\varphi_{\kappa,j}$ from Fig. 3 are formed by the connection of the wires with the same $j$ by common bars through resistances $r_{\kappa,j}$. The potentials $\Phi_{j}$ are applied to the bars. The potentials $\Phi_{j}$ and the resistances $r_{\kappa,j}$ are defined by Eqs. (22) and provide the same currents through the wires as in Fig. 3.
For the considered computational scheme of vector-matrix multiplication based on OQS in each group of close modes, we have $\Phi_{j}(T_{j})=n_{j}(\omega_{\kappa}/T_{j})$ and $r_{\kappa,j}=0$.
As it is shown, electric currents in the circuit presented in Fig. 4 are equivalent to stationary energy flows in the OQS of bosonic modes connected to several reservoirs with different temperatures. The dissipation rates in this OQS are equivalent to resistances at power minus one (conductivities), while reservoirs’ occupancies and temperatures can be seen as potentials (see Figs. 2-4).
Using this analogy, we can expand the class of matrices that can be used in computations similarly to how it is done in MNN. If we need to multiply a matrix $A$ whose rows are normalized vectors with positive and negative elements by a positive vector $\vec{b}$, we can split it into two matrices $A=A_{+}-A_{-}$. The matrix $A_{+}$ consists of all positive elements in their places in the matrix $A$ and zeros in other places. The matrix $A_{-}$ consists of all absolute values of negative elements of $A$ in their places in $A$, and zeros in other places. After that, the computations can be done with both matrices $A_{+}$ and $A_{-}$ separately using Eq. (12). The results being subtracted with attenuation deliver $A\vec{b}$.
Conclusion and discussion. In this letter, we show that OQS of bosonic modes, i.e., photons, phonons, magnons, etc., can be used as thermodynamic coprocessors implementing multiple vector-matrix multiplications in a parallel analog way with stochastic matrices. We show that stationary energy flow through one of the OQS’s modes to a reservoir which is sufficiently colder than others (drain reservoir) is proportional to the scalar product of vectors composed of reservoirs’ occupancies and weighted dissipation rates. We show that this fact and ability to manage dissipation rates of OQS’s modes with close frequencies allow us to implement multiplication of an $n$-dimensional vector by an $m\times n$-dimensional matrix, where $n+1$ is the number of reservoirs and $m$ is the number of the modes having close frequencies. The procedure requires the time needed for the establishment of the stationary energy flows, i.e., the OQS dissipation time, which does not depend on the number of the reservoirs (input vector dimension) and the initial state of the OQS. In parallel, vector-matrix multiplications with different matrices are implemented at other close frequency groups of OQS’s modes.
This computational system can be realized by microring resonators in the GHz range (OQS) and heated waveguides (reservoirs). It is expected to reveal a maximal computational rate of $10\div 1000$ TOps/s on the $5\times 5$ cm2 area per a mode of the OQS. However, with nowadays technological opportunities in temperature manipulation, this computational rate reduces to $100$ GOps/s. Nevertheless, even this sufficiently reduced computational rate is comparable with the computational rate of some modern GPU solutions and is of interest in the context of a full-thermodynamic computer development.
The evolution of the OQS under interaction with the thermal environments is treated in the global approach to the dissipation [47, 93, 127] that ensures the fulfillment of the second law of thermodynamics. Hence, the computations being implemented by energy flows are accompanied by entropy growth due to OQS evolution towards its stationary state. Thus, the presented computational scheme can be seen as a part of developing fields of thermal linear algebra, thermodynamic computing, and thermal neural networks [3, 77, 70, 118].
Notably, intense interaction with the environment in many applications is considered to be a drawback, i.e., entanglement, coherence, and energy should be preserved as long as possible in the OQS  [104, 123, 71]. In the considered computational scheme, the situation is opposite. The faster OQS dissipates, the faster computations are done, particularly, the lower the Q-factor of the microring resonator is, the faster computations are done.
We develop an electrical analogy for the discussed computing system. We show that the system is equivalent to a CS. The dissipation rates of the OQS multiplied by the OQS’s modes’ frequencies can be seen as conductivities, reservoirs’ occupancies as potentials, and stationary energy flows as electric currents.
Also, we note that the similar dependence of energy and particle flows on the weighted occupancy of reservoirs is found in fermionic systems with quadratic Hamiltonians [126], i.e., the same electrical analogy holds in this case. Thus, these systems can also be used for the implementation of vector-matrix multiplication at the nanoscale.
Acknowledgement. A.A.Z. and E.S.A. acknowledge the support of the Foundation for the Advancement of Theoretical Physics and Mathematics BASIS.